In the context of multiple regression, define the n X n matrix M =- X(XX)-X. (i) Show that M is symmetric and idempotent.
Answer #1

solution: Let, Mz In- x(xx)x=001 - M= In-xcxx)x M = [In - x(xx)x) ; In-In den for L (AB) = B A 21 - [* (xx)x): Mis symmetric and idempotent. M = In - x[x)? x , 100 - - .: In = 0 0- -- DOO-- Diagonal Elements of In=1, and We also knowGauss-markov v Y = xß tu es satisfies the Akscembling, ECU)=0 Varlu) = ? In Bors = (xx) xy 11 na Now, û ry- 1 = y = x(xxJ*iv) and under Gause markave assumptions Y=xß tl, E(M) =0, var( ) = In واد(دعا = امام → ols residuals î=y- ý ro O amy Covců)